Four hundred grams of aluminum at 100 ᵒ C is dropped into 100 g of water at 20 ᵒ C

contained in a copper calorimeter of mass 80 g. What is the final temperature of the mixture?


1
Expert's answer
2021-05-10T09:17:43-0400

for aluminum

mass M1=400g

T1=100+273=373K

specific heat of aluminum S1=0.902

for copper

M2=100g

T2=20+273=293K

S2=0.385


M1S1(T1-T)=M2S2(T-T2)

3.608(373-T)=0.385(T-293)

9.3(373-T)=T-293

3468.9-9.3T=T-293

10.3T=3761

T=365K

T=365-273

T=92OC